# Existence of Solutions to the Seiberg-Witten Vortex Equations with Exponential Decay on the Plane

### v1

@@ 1.1. Main Results
The derivation of the Seiberg-Witten vortex equations is inspired by the work of Hitchin. In [MR0887284], Hitchin studied a class of solutions to the (anti)self-dual Yang-Mills equations on $\mathbf{R}^{4}$ that are translational invariant in the last two coordinates. The resulting equations on $\mathbf{R}^{2}$ are now called the Hitchin equations. The Hitchin moduli space on a Riemann surface has many striking properties, one of which is the correspondence with the moduli space of Higgs bundle [hitchin1987stable]. The exploitation of this correspondence led to many applications in number theory, notably the work of Ngô in the Langlands program [ngo2006fibration]. For this reason, it seems interesting and potentially fruitful to apply the same pathway to the Seiberg-Witten equations, as already initiated by Dey’s work [MR1952133, MR2698453]. Note that the dimensional reduction of this type is important in establishing a Topological Quantum Field Theory (TQFT) of dimension $(2+1)$.
@@
$\begin{cases}2\dfrac{\partial\psi_{2}}{\partial\overline{z}}+i(A_{0}+iA_{1})\psi_{2}=0,\\ 2\dfrac{\partial\psi_{1}}{\partial z}+i(A_{0}-iA_{1})\psi_{1}=0,\\ i\left(\dfrac{\partial A_{1}}{\partial x_{0}}-\dfrac{\partial A_{0}}{\partial x_{1}}\right)=\dfrac{i}{2}(|\psi_{1}|^{2}-|\psi_{2}|^{2}).\end{cases}$
Here $A_{0},A_{1}$ are real-valued functions on $\mathbf{R}^{2}$, and $\psi_{1},\psi_{2}$ are complex-valued functions on $\mathbf{R}^{2}$. The Seiberg-Witten vortex equations have a symmetry given by the gauge group $\mathcal{G}=Maps(\mathbf{R}^{2},U(1))$, i.e, the solutions are invariant under the action of the gauge group $\mathcal{G}$. By moduli space of the equations, we mean the space of solutions quotient out by the $\mathcal{G}$-symmetry. It is not difficult to write down a solution of (4.1), albeit "trivial". Here is one: $(\partial f/\partial x_{0},\partial f/\partial x_{1},0,0)$, where $f$ is any smooth function on $\mathbf{R}^{2}$. In fact, one can be a little bit algebraically creative and realize that
@@
For any $(c_{1},c_{2})\in\mathbf{C}^{*}\times\mathbf{C}^{*}$ and $\theta\in\mathbf{R}$, $(A_{0},A_{1},\psi_{1},\psi_{2})$ given by
@@
The heart of the analysis that goes into the proof of Theorem 1.3 is the existence and uniqueness result of a certain Kazdan-Warner equation with an appropriate boundary condition.
@@
The exponential decay in Property (E) seems to be out of place. However, it is quite natural to consider. The Seiberg-Witten vortex equations can be thought of as a variant of the vortex equations which were introduced by Ginzburg and Landau to study the theory of superconductivity. From the Mathematics perspective, they are the absolute minimum condition for the Yang-Mills-Higgs functional (cf. Subsection 3.1). Many authors have studied the existence of solutions related to these types of vortex equations derived from Yang-Mills-Higgs models (see, e.g, [MR0614447, MR0573986, garcia1994direct, bradlow1990vortices, bradlow1991special, MR4516074]). Using the Yang-Mills-Higgs functional, the solutions of the Ginzburg-Landau vortex equations can be shown to have exponential decay a priori. To the best of our knowledge, we do not know whether our (4.1) is also the absolute minimum condition for some Yang-Mills-Higgs-type functional. Regardless, we note that by Theorem 1.1, Theorem 1.2 and Theorem 1.3, solutions of (4.1) can exhibit both polynomial growth or exponential decay, a feature that is not shared with the Ginzburg-Landau vortex equations.
As a final remark, the first two equations of the Seiberg-Witten vortex equations can be thought of as a system of Vekua equations. This insight is crucial for us to study the zeroes of exponential decay solutions of (4.1). We develop the basic theory of solutions to these systems of this type that we require and associate the functions that solve these systems with the solutions of vortex equations. This is the first time that a system of Vekua equations of this kind (to be precise a Vekua equation and the result of applying complex conjugation to both sides of a Vekua equation) has been studied. This connection between the classical complex analysis structure of Vekua equations and the gauge equations of mathematical physics is novel. See [Vek], [Bers], and Section 2 for background on Vekua equations.
@@ Acknowledgement.
Parts of this work were carried out during the first-named author’s visit to Washington University in St. Louis in the Fall of 2023. The authors would like to thank the generous accommodation of the university. The second-named author is also grateful for many conversations with Aliakbar Daemi regarding various vortex equations on Riemann surfaces. ∎
@@ 2.1. Classical Vekua Equation
$\phi(z):=\begin{cases}T\left(A+B\dfrac{\overline{w}}{w}\right)(z),&w(z)\neq 0\\ 0,&w(z)=0\end{cases}.$
@@
$\phi(z):=\begin{cases}T\left(A+B\dfrac{\overline{w}}{w}+\dfrac{f}{w}\right)(z),&w(z)\neq 0\\ 0,&w(z)=0\end{cases}$
@@ 2.2. System of Vekua Equations
$\phi_{1}(z)=\begin{cases}T\left(-\left(\overline{iA+B\dfrac{w_{2}}{w_{1}}}\right)\right)(z),&w_{1}(z)\neq 0\\ 0,&w_{1}(z)=0\end{cases}$
@@
$\phi_{2}(z)=\begin{cases}T\left(-\left(i\overline{A}+\overline{B}\dfrac{w_{1}}{w_{2}}\right)\right)(z),&w_{2}(z)\neq 0\\ 0,&w_{2}(z)=0\end{cases}$
@@
$\varphi_{1}(z):=\begin{cases}T\left(-\left(\overline{iA+B\dfrac{w_{2}}{w_{1}}}\right)\right)(z),&w_{1}(z)\neq 0\\ 0,&w(z)=0\end{cases}$
@@
$\varphi_{2}(z)=\begin{cases}T\left(-\left(i\overline{A}+\overline{B}\dfrac{w_{1}}{w_{2}}\right)\right)(z),&w_{2}(z)\neq 0\\ 0,&w_{2}(z)=0\end{cases}$
@@ 3. Vortex Equations
This section presents a brief introduction to mathematical gauge theory and sets up relevant terminologies that go into the statement of the main result of the paper. For more details, we direct the readers to the following references [MR0614447, MR0573986, MR3012377, MR1367507, MR1339810, MR1306021, MR2698453, MR0887284, SACLIOGLU1997675, MR1362874].
@@ 3.1. Gauge Theory Set-Up
  • A gauge potential $A=\sum_{j=1}^{n}A_{j}(x)dx_{j}$. In differential geometry, gauge potentials are to be understood as connections of certain principal $G$-bundle $P$ over $\mathbf{R}^{n}$, where $G$ is a Lie group. Thus, $A_{j}$ are functions defined on $\mathbf{R}^{n}$ that take values in the Lie algebra $\mathfrak{g}$ of $G$.
@@
Here, $x$ denotes a point in $\mathbf{R}^{n}$, written in terms of the standard coordinate system as $x=(x_{1},\cdots,x_{n})$. Let $\{e_{1},\ldots,e_{n}\}$ be the standard orthonormal basis of $\mathbf{R}^{n}$. $G$ is called the gauge group, i.e, it is the group of transformations of the internal symmetry space $E$ that $\phi$ takes value in. The interaction between a gauge potential and matter field is via the notion of taking a covariant derivative. In particular,
$\nabla_{A}\phi=\displaystyle\sum_{j=1}^{n}(d\phi(e_{j})+\rho(A_{j})(\phi))dx_{j}.$
@@
In what follows, we consider the case where $n=2$, the gauge group $G=U(1)$, $P$ is taken to be the trivial principal bundle $P=\mathbf{R}^{2}\times U(1)$, and $\rho:U(1)\to GL(\mathbf{C})$ to be the standard representation of $U(1)$ on $\mathbf{C}$ given by the (complex) scalar multiplication. As a result, the associated vector bundle $E$ simplifies to be the trivial complex line bundle $\mathbb{L}$. Since we are working on Euclidean space $\mathbf{R}^{2}$ where there is no interesting topology and the Lie algebra of $U(1)$ is simply $i\mathbf{R}$, various actors defined above can be simplified as follows. The matter field $\phi$ is now simply a $\mathbf{C}$-valued function on $\mathbf{R}^{2}$. The gauge potential $A$ is a purely imaginary-valued $1$-form on $\mathbf{R}^{2}$, written as $A=iA_{1}dx_{1}+iA_{2}dx_{2}$. Hence, the curvature $F(A)$ would just be the curl of $A$ given by
$F(A)=i\left(\dfrac{\partial A_{2}}{\partial x_{1}}-\dfrac{\partial A_{1}}{\partial x_{2}}\right)dx_{1}\wedge dx_{2}.$
@@
The various point-wise norms that appear in the integrand above deserve some justification. Firstly, since $\phi$ is complex-valued, $|\phi|^{2}=\phi\overline{\phi}$. Next, since $\nabla_{A}\phi=\sum_{j=1}^{2}(d\phi(e_{j})+iA_{j}\phi)dx_{j}$ is a $\mathbf{C}$-valued $1$-form on $\mathbf{R}^{2}$, we simply define $|\nabla_{A}\phi|^{2}$ to be the sum of the squares of the norm of the $\mathbf{C}$-component of the form. Similarly,
$|F(A)|^{2}=\left|\dfrac{\partial A_{2}}{\partial x_{1}}-\dfrac{\partial A_{1}}{\partial x_{2}}\right|^{2}.$
@@
Let $\phi$ be written as $\phi_{1}+i\phi_{2}$ when we view it as a complex-valued function defined on $\mathbf{R}^{2}$ or $\begin{bmatrix}\phi_{1}&\phi_{2}\end{bmatrix}^{T}$ when viewed as a map from $\mathbf{R}^{2}\to\mathbf{R}^{2}$, where $\phi_{j}$ are $\mathbf{R}$-valued functions. We can re-write the Yang-Mills-Higgs action functional (3.1) as follows
$\mathscr{Y}(A,\phi)=\dfrac{1}{2}\iint_{\mathbf{R}^{2}}$ $[(\partial_{1}\phi_{1}-A_{1}\phi_{2})+(\partial_{2}\phi_{2}+A_{2}\phi_{1})]^{2}+$
$+[(\partial_{1}\phi_{2}+A_{1}\phi_{1})-(\partial_{2}\phi_{1}-A_{2}\phi_{2})]^{2}+$
$+\dfrac{1}{2}\iint_{\mathbf{R}^{2}}((\partial_{1}A_{2}-\partial_{2}A_{1})/2+|\phi|^{2}-1)^{2}+\dfrac{1}{2}\iint_{\mathbf{R}^{2}}F(A)+$
$+\iint_{\mathbf{R}^{2}}d(\phi_{2}(d\phi_{1}-iA\phi_{2})-\phi_{1}(d\phi_{2}+iA\phi_{1})).$
@@
When we view $\phi=\begin{bmatrix}\phi_{1}&\phi_{2}\end{bmatrix}^{T}:\mathbf{R}^{2}\to\mathbf{R}^{2}$, its total derivative in matrix form is written as
$d\phi=\begin{bmatrix}\partial_{1}\phi_{1}&\partial_{2}\phi_{1}\\ \partial_{1}\phi_{2}&\partial_{2}\phi_{2}\end{bmatrix}.$
As a result, $d\phi(e_{1})=\partial_{1}\phi_{1}+i\partial_{1}\phi_{2}$ and $d\phi(e_{2})=\partial_{2}\phi_{1}+i\partial_{2}\phi_{2}$. Thus, from the definition of the Yang-Mills-Higgs functional, we can re-write
$\mathscr{Y}(A,\phi)=\dfrac{1}{2}\iint_{\mathbf{R}^{2}}$ $\dfrac{1}{4}(\partial_{1}A_{2}-\partial_{2}A_{1})^{2}+|(\partial_{1}\phi_{1}-A_{1}\phi_{2})+i(\partial_{1}\phi_{2}+A_{1}\phi_{1})|^{2}+$
$+|(\partial_{2}\phi_{1}-A_{2}\phi_{2})+i(\partial_{2}\phi_{2}+A_{2}\phi_{1})|^{2}+(|\phi|^{2}-1)^{2}.$
@@
$\mathscr{Y}(A,\phi)=$ $\dfrac{1}{2}\iint_{\mathbf{R}^{2}}[(\partial_{1}\phi_{1}-A_{1}\phi_{2})+(\partial_{2}\phi_{2}+A_{2}\phi_{1})]^{2}+$
$+\dfrac{1}{2}\iint_{\mathbf{R}^{2}}[(\partial_{1}\phi_{2}+A_{1}\phi_{1})-(\partial_{2}\phi_{1}-A_{2}\phi_{2})]^{2}+$
$+\dfrac{1}{2}\iint_{\mathbf{R}^{2}}((\partial_{1}A_{2}-\partial_{2}A_{1})/2+|\phi|^{2}-1)^{2}+\dfrac{1}{2}\iint_{\mathbf{R}^{2}}F(A)+$
$+\dfrac{1}{2}\iint_{\mathbf{R}^{2}}-(\partial_{1}A_{2}-\partial_{2}A_{1})|\phi|^{2}+$
$+\dfrac{1}{2}\iint_{\mathbf{R}^{2}}-2(\partial_{1}\phi_{1}-A_{1}\phi_{2})(\partial_{2}\phi_{2}+A_{2}\phi_{1})+$
$+\dfrac{1}{2}\iint_{\mathbf{R}^{2}}2(\partial_{1}\phi_{2}+A_{1}\phi_{1})(\partial_{2}\phi_{1}-A_{2}\phi_{2}).$
@@
$-(\partial_{1}A_{2}-\partial_{2}A_{1})|\phi|^{2}-2(\partial_{1}\phi_{1}-A_{1}\phi_{2})(\partial_{2}\phi_{2}+A_{2}\phi_{1})+2(\partial_{1}\phi_{2}+A_{1}\phi_{1})(\partial_{2}\phi_{1}-A_{2}\phi_{2})$
$=\underbrace{(-\partial_{1}A_{2}\cdot\phi^{2}_{1}-2A_{2}\phi_{2}\cdot\partial_{1}\phi_{1})}_{\text{$-\partial_{1}(A_{2}\phi^{2}_{1})$}}+\underbrace{(-\partial_{1}A_{2}\cdot\phi^{2}_{2}-2A_{2}\phi_{2}\cdot\partial_{1}\phi_{2})}_{\text{$-\partial_{1}(A_{2}\phi^{2}_{2})$}}+$
$+\underbrace{(\partial_{2}A_{1}\cdot\phi^{2}_{1}+2A_{1}\phi_{1}\cdot\partial_{1}\phi_{2})}_{\text{$\partial_{2}(A_{1}\phi^{2}_{1})$}}+\underbrace{(\partial_{2}A_{1}\cdot\phi^{2}_{2}+2A_{1}\phi_{2}\cdot\partial_{2}\phi_{2})}_{\text{$\partial_{2}(A_{1}\phi^{2}_{2})$}}+$
$+\underbrace{2(-\partial_{1}\phi_{1}\cdot\partial_{2}\phi_{2}+\partial_{1}\phi_{2}\cdot\partial_{2}\phi_{1})}_{\text{$-2\det(d\phi)$}}.$
@@
$d(\phi_{1}d\phi_{2}-\phi_{2}d\phi_{1})$ $=2d\phi_{1}\wedge d\phi_{2}=2\det(d\phi)dx_{1}\wedge dx_{2},$
$d(iA|\phi|^{2})=-d(A_{2}|\phi|^{2}dx_{2}+A_{1}|\phi|^{2}dx_{1})$ $=(-\partial_{1}(A_{2}|\phi|^{2})+\partial_{2}(A_{1}|\phi|^{2}))dx_{1}\wedge dx_{2}.$
@@
(3.2) $\begin{cases}\dfrac{\partial\phi}{\partial\overline{z}}-\dfrac{i}{2}(A_{1}+iA_{2})\phi=0,\\ \dfrac{i}{2}F(A)=(1-|\phi|^{2})dz\wedge d\overline{z},\end{cases}$
where $A=iA_{1}dx_{1}+iA_{2}dx_{2}$.
@@
Note that $\phi_{2}(d\phi_{1}-iA\phi_{2})-\phi_{1}(d\phi_{2}+iA\phi_{1})$ can be re-written as
$\left\langle\begin{bmatrix}d&-iA\\ iA&d\end{bmatrix}\,\begin{bmatrix}\phi_{1}\\ \phi_{2}\end{bmatrix},\begin{bmatrix}\phi_{2}\\ -\phi_{1}\end{bmatrix}\right\rangle=\Re\langle\nabla_{A}\phi,-i\phi\rangle,$

### v2

@@ 1.1. Main Results
The derivation of the Seiberg-Witten vortex equations is inspired by the work of Hitchin. In [MR0887284], Hitchin studied a class of solutions to the (anti)self-dual Yang-Mills equations on $\mathbf{R}^{4}$ that are translational invariant in the last two coordinates. The resulting equations on $\mathbf{R}^{2}$ are now called the Hitchin equations. The Hitchin moduli space on a Riemann surface has many striking properties, one of which is the correspondence with the moduli space of Higgs bundle [hitchin1987stable]. The exploitation of this correspondence led to many applications in number theory, notably the work of Ngô in the Langlands program [ngo2006fibration]. For this reason, it seems interesting and potentially fruitful to apply the same pathway to the Seiberg-Witten equations, as already initiated by Dey’s work [MR1952133, MR2698453].
@@
(1.1) $\begin{cases}2\dfrac{\partial\psi_{2}}{\partial\overline{z}}+i(A_{0}+iA_{1})\psi_{2}=0,\\ 2\dfrac{\partial\psi_{1}}{\partial z}+i(A_{0}-iA_{1})\psi_{1}=0,\\ i\left(\dfrac{\partial A_{1}}{\partial x_{0}}-\dfrac{\partial A_{0}}{\partial x_{1}}\right)=\dfrac{i}{2}(|\psi_{1}|^{2}-|\psi_{2}|^{2}),\end{cases}$
where $z=x_{0}+ix_{i}$, $x_{0}$ and $x_{1}$ are the standard real coordinates, and we identify $\mathbf{C}$ with $\mathbf{R}^{2}$. Here $A_{0},A_{1}$ are real-valued functions on $\mathbf{R}^{2}$, and $\psi_{1},\psi_{2}$ are complex-valued functions on $\mathbf{R}^{2}$. The Seiberg-Witten vortex equations have a symmetry given by the gauge group $\mathcal{G}=Maps(\mathbf{R}^{2},U(1))$, i.e, the solutions are invariant under the action of the gauge group $\mathcal{G}$. By moduli space of the equations, we mean the space of solutions quotient out by the $\mathcal{G}$-symmetry. It is not difficult to write down a solution of (1.1), albeit "trivial". Here is one: $(\partial f/\partial x_{0},\partial f/\partial x_{1},0,0)$, where $f$ is any smooth function on $\mathbf{R}^{2}$. In fact, one can be a little bit algebraically creative and realize that
@@
Let $\{z_{1},\cdots,z_{k}\}$ be a finite collection of points on the plane. For any $(c_{1},c_{2})\in\mathbf{C}^{*}\times\mathbf{C}^{*}$ and $\theta\in\mathbf{R}$, $(A_{0},A_{1},\psi_{1},\psi_{2})$ given by
@@
Definition 1.3.
A solution $(A_{0},A_{1},\psi_{1},\psi_{2})$ of (1.1) is said to satisfy property (E) if and only if $\psi_{1},\psi_{2}$ have the same argument and that for every $\epsilon>0$, there exists a $c>0$ such that $C-|\psi_{2}|^{2}\leq c\cdot\exp(-(1-\epsilon)|x|)$ for some fixed constant $C$.
@@
Let $C$ be any positive number. There exists a smooth solution $(A_{0},A_{1},\psi_{1},\psi_{2})$ of (1.1) with the property (E).
The heart of the analysis that goes into the proof of Theorem 1.4 is the existence result of a certain Kazdan-Warner equation with an appropriate boundary condition. The following theorem might be well-known to experts, but we could not find a direct reference of it in the literature. In a later section (cf. Section 4), we give a detailed proof of the theorem.
@@
The exponential decay in Property (E) seems to be out of place. However, it is quite natural to consider. The Seiberg-Witten vortex equations can be thought of as a variant of the vortex equations which were introduced by Ginzburg and Landau to study the theory of superconductivity. From the Mathematics perspective, they are the absolute minimum condition for the Yang-Mills-Higgs functional (cf. Subsection 3.1). Many authors have studied the existence of solutions related to these types of vortex equations derived from Yang-Mills-Higgs models (see, e.g, [MR0614447, MR0573986, garcia1994direct, bradlow1990vortices, bradlow1991special, MR4516074, MR3435967, MR3513572]). We note that in [MR3435967, MR3513572], the authors’ main results are generalizations of Taubes’ result [MR0573986, MR0614447] for various gauged-sigma models of the Ginzburg-Landau vortex equations. Using the Yang-Mills-Higgs functional, the solutions of the Ginzburg-Landau vortex equations can be shown to have exponential decay a priori. To the best of our knowledge, we do not know whether our equation (1.1) is also the absolute minimum condition for some Yang-Mills-Higgs-type functional. Regardless, we note that by Theorem 1.1, Theorem 1.2 and Theorem 1.4, solutions of (1.1) can exhibit both polynomial growth or exponential decay, a feature that is not shared with the Ginzburg-Landau vortex equations.
As a final remark, the first two equations of the Seiberg-Witten vortex equations can be thought of as a system of Vekua-type equations. The classical Vekua equation is the canonical form of a first-order elliptic equation in the plane and generalizes the classical Cauchy-Riemann equation from complex analysis, which appears naturally in the study of Ginzburg-Landau vortex equations. The connection between Seiberg-Witten vortex equations and Vekua equations is crucial for us to study the zeroes of exponential decay solutions of (1.1). We develop the basic theory of solutions to the systems of Vekua-type equations that we require and associate the functions that solve these systems with the solutions of vortex equations. This is the first time that a system of Vekua equations of this kind (to be precise a Vekua equation and the result of applying complex conjugation to both sides of a Vekua equation) has been studied. This connection between the classical complex analysis structure of Vekua equations and the gauge equations of mathematical physics is novel. See [Vek, Bers], and Section 2 for background on Vekua equations.
@@ Acknowledgement.
Parts of this work were carried out during the first-named author’s visit to Washington University in St. Louis in the Fall of 2023. The authors would like to thank the university for their generous accommodations. The second-named author is also grateful for many conversations with Aliakbar Daemi regarding various vortex equations on Riemann surfaces. ∎
@@ 2.1. Classical Vekua Equation
$\phi(z):=\begin{cases}T\left(A+B\dfrac{\overline{w}}{w}\right)(z),&w(z)\neq 0\\ T\left(A+B\right)(z),&w(z)=0\end{cases}.$
@@
$\phi(z):=\begin{cases}T\left(A+B\dfrac{\overline{w}}{w}+\dfrac{f}{w}\right)(z),&w(z)\neq 0\\ T\left(A+B+\dfrac{f}{w}\right)(z),&w(z)=0\end{cases}$
@@ 2.2. System of Vekua Equations
$\phi_{1}(z)=T\left(-\left(\overline{iA+B\dfrac{w_{2}}{w_{1}}}\right)\right)(z)$
@@
$\phi_{2}(z)=T\left(-\left(i\overline{A}+\overline{B}\dfrac{w_{1}}{w_{2}}\right)\right)(z)$
@@
$\phi_{1}(z)=T\left(-\left(\overline{iA+B\dfrac{w_{2}}{w_{1}}}\right)\right)(z)$
@@
$\phi_{2}(z)=T\left(-\left(i\overline{A}+\overline{B}\dfrac{w_{1}}{w_{2}}\right)\right)(z)$
@@ 3. Vortex Equations
This section presents a brief introduction to mathematical gauge theory and sets up relevant terminologies that go into the statement of the main result of the paper. For more details, we direct the readers to the following references [MR0614447, MR0573986, MR3012377, MR1367507, MR1339810, MR1306021, MR2698453, MR0887284, SACLIOGLU1997675, MR1362874]. The material in this section is well-known to experts. Our intention for providing detailed proofs of these included results is for self-containment and accessibility to a larger audience.
@@ 3.1. Gauge Theory Set-Up
  • A gauge potential $A=\sum_{j=0}^{n-1}A_{j}(x)dx_{j}$. In differential geometry, gauge potentials are to be understood as connections of certain principal $G$-bundle $P$ over $\mathbf{R}^{n}$, where $G$ is a Lie group. Thus, $A_{j}$ are functions defined on $\mathbf{R}^{n}$ that take values in the Lie algebra $\mathfrak{g}$ of $G$.
@@
Here, $x$ denotes a point in $\mathbf{R}^{n}$, written in terms of the standard coordinate system as $x=(x_{0},\cdots,x_{n-1})$. Let $\{e_{0},\ldots,e_{n-1}\}$ be the standard orthonormal basis of $\mathbf{R}^{n}$. $G$ is called the gauge group, i.e, it is the group of transformations of the internal symmetry space $E$ that $\phi$ takes value in. The interaction between a gauge potential and matter field is via the notion of taking a covariant derivative. In particular,
$\nabla_{A}\phi=\displaystyle\sum_{j=0}^{n-1}(d\phi(e_{j})+\rho(A_{j})(\phi))dx_{j}.$
@@
In what follows, we consider the case where $n=2$, the gauge group $G=U(1)$, $P$ is taken to be the trivial principal bundle $P=\mathbf{R}^{2}\times U(1)$, and $\rho:U(1)\to GL(\mathbf{C})$ to be the standard representation of $U(1)$ on $\mathbf{C}$ given by the (complex) scalar multiplication. As a result, the associated vector bundle $E$ simplifies to be the trivial complex line bundle $\mathbb{L}$. Since we are working on Euclidean space $\mathbf{R}^{2}$ where there is no interesting topology and the Lie algebra of $U(1)$ is simply $i\mathbf{R}$, various actors defined above can be simplified as follows. The matter field $\phi$ is now simply a $\mathbf{C}$-valued function on $\mathbf{R}^{2}$. The gauge potential $A$ is a purely imaginary-valued $1$-form on $\mathbf{R}^{2}$, written as $A=iA_{0}dx_{0}+iA_{1}dx_{1}$. Hence, the curvature $F(A)$ would just be the curl of $A$ given by
$F(A)=i\left(\dfrac{\partial A_{1}}{\partial x_{0}}-\dfrac{\partial A_{0}}{\partial x_{1}}\right)dx_{0}\wedge dx_{1}.$
@@
The various point-wise norms that appear in the integrand above deserve some justification. Firstly, since $\phi$ is complex-valued, $|\phi|^{2}=\phi\overline{\phi}$. Next, since $\nabla_{A}\phi=\sum_{j=0}^{1}(d\phi(e_{j})+iA_{j}\phi)dx_{j}$ is a $\mathbf{C}$-valued $1$-form on $\mathbf{R}^{2}$, we simply define $|\nabla_{A}\phi|^{2}$ to be the sum of the squares of the norm of the $\mathbf{C}$-component of the form. Similarly,
$|F(A)|^{2}=\left|\dfrac{\partial A_{1}}{\partial x_{0}}-\dfrac{\partial A_{0}}{\partial x_{1}}\right|^{2}.$
@@
Let $\phi$ be written as $\phi_{0}+i\phi_{1}$ when we view it as a complex-valued function defined on $\mathbf{R}^{2}$ or $\begin{bmatrix}\phi_{0}&\phi_{1}\end{bmatrix}^{T}$ when viewed as a map from $\mathbf{R}^{2}\to\mathbf{R}^{2}$, where $\phi_{j}$ are $\mathbf{R}$-valued functions. We can re-write the Yang-Mills-Higgs action functional (3.1) as follows
$\mathscr{Y}(A,\phi)=\dfrac{1}{2}\iint_{\mathbf{R}^{2}}$ $[(\partial_{0}\phi_{0}-A_{0}\phi_{1})+(\partial_{1}\phi_{1}+A_{1}\phi_{0})]^{2}+$
$+[(\partial_{0}\phi_{1}+A_{0}\phi_{0})-(\partial_{1}\phi_{0}-A_{1}\phi_{1})]^{2}+$
$+\dfrac{1}{2}\iint_{\mathbf{R}^{2}}((\partial_{0}A_{1}-\partial_{1}A_{0})/2+|\phi|^{2}-1)^{2}+\dfrac{1}{2}\iint_{\mathbf{R}^{2}}F(A)+$
$+\iint_{\mathbf{R}^{2}}d(\phi_{1}(d\phi_{0}-iA\phi_{1})-\phi_{0}(d\phi_{1}+iA\phi_{0})).$
@@
When we view $\phi=\begin{bmatrix}\phi_{0}&\phi_{1}\end{bmatrix}^{T}:\mathbf{R}^{2}\to\mathbf{R}^{2}$, its total derivative in matrix form is written as
$d\phi=\begin{bmatrix}\partial_{0}\phi_{0}&\partial_{1}\phi_{0}\\ \partial_{0}\phi_{1}&\partial_{1}\phi_{1}\end{bmatrix}.$
As a result, $d\phi(e_{0})=\partial_{0}\phi_{0}+i\partial_{0}\phi_{1}$ and $d\phi(e_{1})=\partial_{1}\phi_{0}+i\partial_{1}\phi_{1}$. Thus, from the definition of the Yang-Mills-Higgs functional, we can re-write
$\mathscr{Y}(A,\phi)=\dfrac{1}{2}\iint_{\mathbf{R}^{2}}$ $\dfrac{1}{4}(\partial_{0}A_{1}-\partial_{1}A_{0})^{2}+|(\partial_{0}\phi_{0}-A_{0}\phi_{1})+i(\partial_{0}\phi_{1}+A_{0}\phi_{0})|^{2}+$
$+|(\partial_{1}\phi_{0}-A_{1}\phi_{1})+i(\partial_{1}\phi_{1}+A_{1}\phi_{0})|^{2}+(|\phi|^{2}-1)^{2}.$
@@
$\mathscr{Y}(A,\phi)=$ $\dfrac{1}{2}\iint_{\mathbf{R}^{2}}[(\partial_{0}\phi_{0}-A_{0}\phi_{1})+(\partial_{1}\phi_{1}+A_{1}\phi_{0})]^{2}+$
$+\dfrac{1}{2}\iint_{\mathbf{R}^{2}}[(\partial_{0}\phi_{1}+A_{0}\phi_{0})-(\partial_{1}\phi_{0}-A_{1}\phi_{1})]^{2}+$
$+\dfrac{1}{2}\iint_{\mathbf{R}^{2}}((\partial_{0}A_{1}-\partial_{1}A_{0})/2+|\phi|^{2}-1)^{2}+\dfrac{1}{2}\iint_{\mathbf{R}^{2}}F(A)+$
$+\dfrac{1}{2}\iint_{\mathbf{R}^{2}}-(\partial_{0}A_{1}-\partial_{1}A_{0})|\phi|^{2}+$
$+\dfrac{1}{2}\iint_{\mathbf{R}^{2}}-2(\partial_{0}\phi_{0}-A_{0}\phi_{1})(\partial_{1}\phi_{1}+A_{1}\phi_{0})+$
$+\dfrac{1}{2}\iint_{\mathbf{R}^{2}}2(\partial_{0}\phi_{1}+A_{0}\phi_{0})(\partial_{1}\phi_{0}-A_{1}\phi_{1}).$
@@
$-(\partial_{0}A_{1}-\partial_{1}A_{0})|\phi|^{2}-2(\partial_{0}\phi_{0}-A_{0}\phi_{1})(\partial_{1}\phi_{1}+A_{1}\phi_{0})+2(\partial_{0}\phi_{1}+A_{0}\phi_{0})(\partial_{1}\phi_{0}-A_{1}\phi_{1})$
$=\underbrace{(-\partial_{0}A_{1}\cdot\phi^{2}_{0}-2A_{1}\phi_{1}\cdot\partial_{0}\phi_{0})}_{\text{$-\partial_{0}(A_{1}\phi^{2}_{0})$}}+\underbrace{(-\partial_{0}A_{1}\cdot\phi^{2}_{1}-2A_{1}\phi_{1}\cdot\partial_{0}\phi_{1})}_{\text{$-\partial_{0}(A_{1}\phi^{2}_{1})$}}+$
$+\underbrace{(\partial_{1}A_{0}\cdot\phi^{2}_{0}+2A_{0}\phi_{0}\cdot\partial_{0}\phi_{1})}_{\text{$\partial_{1}(A_{0}\phi^{2}_{0})$}}+\underbrace{(\partial_{1}A_{0}\cdot\phi^{2}_{1}+2A_{0}\phi_{1}\cdot\partial_{1}\phi_{1})}_{\text{$\partial_{1}(A_{0}\phi^{2}_{1})$}}+$
$+\underbrace{2(-\partial_{0}\phi_{0}\cdot\partial_{1}\phi_{1}+\partial_{0}\phi_{1}\cdot\partial_{1}\phi_{0})}_{\text{$-2\det(d\phi)$}}.$
@@
$d(\phi_{0}d\phi_{1}-\phi_{1}d\phi_{0})$ $=2d\phi_{0}\wedge d\phi_{1}=2\det(d\phi)dx_{0}\wedge dx_{1},$
$d(iA|\phi|^{2})=-d(A_{1}|\phi|^{2}dx_{1}+A_{0}|\phi|^{2}dx_{0})$ $=(-\partial_{0}(A_{1}|\phi|^{2})+\partial_{1}(A_{0}|\phi|^{2}))dx_{0}\wedge dx_{1}.$
@@
(3.2) $\begin{cases}\dfrac{\partial\phi}{\partial\overline{z}}-\dfrac{i}{2}(A_{0}+iA_{1})\phi=0,\\ \dfrac{i}{2}F(A)=(1-|\phi|^{2})dz\wedge d\overline{z},\end{cases}$
where $A=iA_{0}dx_{0}+iA_{1}dx_{1}$.
@@
Note that $\phi_{1}(d\phi_{0}-iA\phi_{1})-\phi_{0}(d\phi_{1}+iA\phi_{0})$ can be re-written as
$\left\langle\begin{bmatrix}d&-iA\\ iA&d\end{bmatrix}\,\begin{bmatrix}\phi_{0}\\ \phi_{1}\end{bmatrix},\begin{bmatrix}\phi_{1}\\ -\phi_{0}\end{bmatrix}\right\rangle=\Re\langle\nabla_{A}\phi,-i\phi\rangle,$